A circuit is constructed with four resistors, one inductor, one battery and a switch as shown. The values for the resistors are: R1-R2-76 Ω, R3-105 Ω and R4-810. The inductance is L-478 mH and the battery voltage is V-24 V 1) The switch has been open for a long time when at time t - o, the switch is closed. What is 11(0), the magnitude of the current through the resistor R1 just after the switch is closed? A uIii 2) What is l), the magnitude of the current that flows through the resistor R1 a very long time after the switch has been closed? A unii 3) What is VL(0), the magnitude of the voltage across the inductor just after the switch is closed? 4) What is lL(-),the magnitude of the current through the inductor after the switch has been closed for a very long time? A uin
Answer #1

1)

Current, I = 24/(76 + 76 + 105 + 81) = 0.071 A

2)

Current, I = 24/(76 + 81) = 0.153 A

3)

Voltage, V = 0.071(76 + 105) = 12.85 V

4)

Current, I = 0.153 A

5)

Current, I = 0.071 A

Comment in case any doubt please rate my answer .....